solve the problem. x=7y+7
2x-6y=12

Answer by Mathpassionate(25)

solve the problem. x=7y+7
2x-6y=12

We substitute x=7y+7 into 2x-6y=12

2x-6y=12
2*(7y+7)-6y=12
(2*7y)+(2*7)-6y=12
14y+14-6y=12
14y-6y=12-14
8y=-2
y=-2/8
y=-1/4

Now, substitute y = -1/4 into x=7y+7
x=7y+7
x=7*(-1/4)+7
x = -7/4 + 7
x = -7/4 + 28/4 (because 28/4 is the same as 7, but we need common denominators (4)
x = (-7 + 28)/4
x = (21)/4
x = 21/4

Then, the answers are:

y=-1/4
x = 21/4
Any questions, please ask me.